Lins garden has an area of 78 square yards. The length of the garden is 5 less than three times its width. What are the dimensions of Lins garden.

Answer by checkley79(3341)

L=3W-5
AREA=LW
78=(3W-5)W
78=3W^2-5W
3W^2-5W-78=0
(3W+13)(W-6)=0
W-6 =0
W=6 ANS FOR THE WIDTH.
L=3*6-5
L=18-5=13 ANS. FOR THE LENGTH.
PROOF:
78=6*13
78=78